HELP Please. I don't have a clue about geometry. Can someone show me how? I missed this question on a quiz.
3. (6 pts) (Using ∆ABC, find the following:
a) The length of side AC
b) The perimeter of ∆ABC
c) The area of ∆ABC
The triangle is a right angle with the length of base (CB) at 12 inches and the hypoteneuse (AB) is 13 inches.
Does the question make sense?




















Answer by Alan3354(21581)

3. (6 pts) (Using ∆ABC, find the following:
a) The length of side AC
b) The perimeter of ∆ABC
c) The area of ∆ABC
The triangle is a right angle with the length of base (CB) at 12 inches and the hypoteneuse (AB) is 13 inches.
---------------------
Use the Pythagorean Theorem:
hyp^2 = AC^2 + BC^2
13^2 = 12^2 + AC^2
AC^2 = 169 - 144 = 25
AC = 5 inches
--------------
Perimeter = 5 + 12 + 13 = 30 inches
-----------
Area = bh/2 = 5*12/2
Area = 30 square inches